In 10 years, Samuel will be 1/2 as old as Roman. Five years ago, Samuel was only 1/3 as old as Roman. How old is Samuel now?
Answer by richwmiller(17219)

(s+10)=(r+10)/2
(s-5)=(r-5)/3
r=50 s=20